# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v2

@@ 4.1 Reward Regret and Constraint Violation of a Pessimistic Policy
Combining the bounds of all terms, and after rearrangement, we get the bound for the Lagrangian $\mathcal{L}(\lambda)$. We can obtain the regret bound for reward or constraint violation through setting $\bf\lambda=0$ or $\lambda_{i}=\rho e_{i}\mathbf{1}_{\{[\sum_{t=1}^{T^{\prime}}y_{i}^{t}]_{+}>0\}}$ for each constraint $i$.
4.2 Reward Regret and Constraint Violation of an Optimistic Policy
When a certain degree of risk is acceptable in exchange for faster learning and higher reward potential (e.g., exploration-heavy tasks or loosely constrained resource allocation), an optimistic threshold can be adopted in Algorithm 2, by setting $\alpha_{\diamond}^{t}=\underline{\alpha}^{t}$ at every episode $t$. Similarly, we prove reward regret and constraint violation of an optimistic policy as below.
@@ 4.2 Reward Regret and Constraint Violation of an Optimistic Policy
Under the same conditions as Theorem 2, with probability at least $1-\delta$, Algorithm 2 under optimistic thresholds guarantees the following regret bounds:
@@
$+\tilde{\mathcal{O}}\left(\sqrt{mSAH^{4}T/\gamma}+\sqrt{m}SAH^{2}\right).$
where $\rho^{\prime}$ is the Slater gap under optimistic thresholds.
This theorem shows that, even with an exploration-heavy strategy, both reward regret and constraint violation remain sublinear over $T$ episodes.
@@
The proof idea is similar to its pessimistic counterpart. We highlight the most significant differences below, given the space constraints. The detailed proof is given in Appendix H.
@@
We can derive a decomposition similar to Equation (5); however, unlike the pessimistic policy case, an additional term $\sum_{i=1}^{T^{\prime}}(\alpha^{t}-\underline{\alpha}^{t})\lambda$ arises, representing the total estimation error across all thresholds in the optimistic scenario. To derive the bound for this term, we invoke arguments analogous to those developed in Lemmas 3 and 4.
@@
Optimistic policies also exhibit distinct behaviour, as shown below.
Lemma 5 (Policy optimality of optimistic policies).
Conditioned on the success event, for any episode $t\in[T^{\prime}]$, under optimistic thresholds, the following holds:
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\leq\tilde{\mathcal{O}}\left((1+m\rho^{\prime})H^{2}\sqrt{T}\right).$
@@ 4.3 Safety-Performance Trade-Off
Our theoretical analysis reveals a fundamental safety-performance trade-off: policies that prioritize strict constraint adherence can incur larger reward regret, while those that aggressively maximise rewards are more prone to constraint violations. Concretely, Theorems 2 and 3 demonstrate that, compared to fixed-threshold baselines, pessimistic (strict) thresholds yield lower violations but higher reward regret, whereas optimistic thresholds exhibit the opposite behaviour.
@@
One may wonder whether it is possible to tune a hyperparameter to leverage this trade-off without resorting to two extreme cases – pessimistic or optimistic. To this end, we introduce a hyperparameter $\tau\in[0,1]$ to interpolate between these extremes. In particular, instead of setting $\alpha_{\diamond}^{t}=\overline{\alpha}^{t}$ or $\underline{\alpha}^{t}$ at episode $t$, we define a blended threshold as follows:
$\alpha_{\tau}^{t}=\tau\underline{\alpha}^{t}+(1-\tau)\overline{\alpha}^{t}.$
Varying $\tau$ adjusts how ‘strict’ or ‘relaxed’ the thresholds are, thereby controlling the balance between safety and performance. While our theorems currently focus on the two boundary cases $\tau=0$ (pessimistic) and $\tau=1$ (optimistic), it is straightforward to extend to an analogous analysis for intermediate $\tau$ values.
@@ 5 Conclusion
In this paper, we propose a novel primal-dual scheme for CMDPs with stochastic thresholds, which utilises samples from environmental interactions to estimate these thresholds. Building on this framework, we introduce SPOT, a primal-dual algorithm for episodic CMDPs with stochastic thresholds that is designed to balance regret minimisation with safety guarantees in an uncertain environment. Furthermore, we leverage either pessimistic or optimistic thresholds in SPOT, respectively. We theoretically prove that SPOT, under both threshold settings, achieves a regret bound of $\tilde{\mathcal{O}}(\sqrt{T})$ and a constraint violation of $\tilde{\mathcal{O}}(\sqrt{T})$ over $T$ episodes, which are in the same order as that of an approach with certain thresholds. Our findings reveal a subtle dance: pessimistic thresholds enforce stricter safety constraints that effectively reduce violations, albeit at the cost of increased reward regret; in the optimistic setting, a reverse situation is observed. This creates a ‘margin’ between safety and performance without relying on prior threshold information. To our best knowledge, SPOT is the first primal-dual algorithm to provide theoretical performance guarantees even when safety thresholds are uncertain, offering insight for exploring real-world applications in safety-critical domains such as autonomous driving and healthcare.
Applicability
This paper primarily focuses on model-based approaches, specifically in the context of episodic CMDPs. While deep reinforcement learning is highly promising, it falls outside the scope of this work. Nonetheless, our work lays the key theoretical foundations for ensuring AI safety in uncertain environments.
@@ Appendix A Summary of Notation
Action space $\mathcal{A}$, with cardinality $A$
@@
Slater gap under pessimistic thresholds $\rho\;=\;(r^{T}q^{\pi^{\star}}(p)-r^{T}q^{\pi_{1}})/\min_{i\in[m]}\xi_{i}$
Slater gap under optimistic thresholds $\rho^{\prime}\;=\;(r^{T}q^{\pi^{\star}}(p)-r^{T}q^{\pi_{2}})/\min_{i\in[m]}\xi^{\prime}_{i}$
@@
Threshold $\tilde{\alpha}_{i,h}(s,a)$ with $\mathbb{E}[\tilde{\alpha}_{i,h}(s,a)]=\alpha_{i,h}(s,a)$
Constraint function $y_{t}(s,a)=\alpha^{t}-Gq^{\pi_{t}}(p)$
@@
Occupancy measure (matrix form) $q^{\pi}(p)=\{q_{h}^{\pi}(s,a;p)\}_{H\times S\times A}$
Lagrangian under stochastic thresholds $\mathcal{L}(q,\lambda,\omega)=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\alpha(\omega))$
Lagrangian of Algorithm 2 $\mathcal{L^{\prime}}(q,\lambda,\omega)=\overline{r}^{T}q^{\pi}(p)+\lambda^{T}(\overline{G}^{T}q^{\pi}(p)-\alpha_{\diamond}(\omega))$
Optimal policy $q^{\star}\in\arg\max_{q\in\mathcal{F}_{q}}\min_{\lambda\in\mathbb{R}_{\geq 0}^{m}}\mathcal{L}(q,\lambda,\omega)$
@@
Stepsize $\eta_{t}$, $\eta_{\lambda}>0$ (hyperparameter)
@@
Visitation counter $N_{h}^{t}(s,a):=\sum_{l=\max\{1,t-W_{t}+1\}}^{t}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$
@@
Optimistic estimates $\overline{\alpha}^{t},\;\overline{r}^{t},\;\overline{g}^{t},\;\overline{p}^{t}$
@@ B.1 Lagrangian Formulation of CMDPs without Stochastic Thresholds
Under the bandit feedback paradigm, traditional CMDPs usually assume that rewards and constraints are stochastic or adversarial. Here, we take the stochastic case as an example, that is, rewards and constraints are randomly drawn from some probability distributions $\mathcal{R}$ and $\mathcal{G}_{i}$ for all $i\in[m]$. The non-stationary immediate reward after taking an action $a$ at state $s$ is a random variable $\tilde{r}_{h}(s,a)\in[0,1]$ with expectation $\mathbb{E}_{\mathcal{R}}[\tilde{r}_{h}(s,a)]=r_{h}(s,a)$. Similarly, for every constraint $i\in[m]$, the non-stationary immediate constraint $\tilde{g}_{i,h}(s,a)\in[0,1]$ for every state-action pair $(s,a)$ is a random variable with expectation $\mathbb{E}_{\mathcal{G}_{i}}[\tilde{g}_{i,h}(s,a)]=g_{i,h}(s,a)$.
@@
$\max_{q\in\Delta(\mathcal{M})}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}(q,\lambda):=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\alpha),$
@@
In order to further discuss its strong duality, we make the following assumption, which is commonly used in the study of CMDPs [1, 11, 18, 40, 9, 10, 22, 26].
Assumption 1 (Slater condition under fixed thresholds).
Given a constraint matrix $G=[g_{1},\ldots,g_{m}]$, there exists a strictly feasible solution $q^{\overline{\pi}}\in\mathcal{F}_{q}$ such that $g_{i}^{T}q^{\overline{\pi}}<\alpha_{i}$ for every constraint $i\in[m]$.
@@
Lemma 6 (Strong duality under fixed thresholds [27]).
@@ B.2 Lagrangian Formulation of CMDPs with Stochastic Thresholds
From these definitions, it follows that the pessimistic and optimistic scenarios define the boundaries for all feasible policies, yielding: $\mathcal{F}_{\textup{pes}}\subseteq\mathcal{F}_{q}(\omega)\subseteq\mathcal{F}_{\textup{opt}}$ with probability at least $1-\delta$ due to Theorem 1.
To identify the corresponding optimal policies within these boundary scenarios, we solve the following constrained optimisation problems:
$\max_{q\in\mathcal{F}_{\textup{pes}}}\min_{\lambda\geq 0}\mathcal{L}(q,\lambda,\omega):=r^{T}q^{\pi}(p)+\lambda^{T}\left(G^{T}q^{\pi}(p)-\overline{\alpha}(\omega)\right),\quad$ (Pessimistic)
$\max_{q\in\mathcal{F}_{\textup{opt}}}\min_{\lambda\geq 0}\mathcal{L}(q,\lambda,\omega):=r^{T}q^{\pi}(p)+\lambda^{T}\left(G^{T}q^{\pi}(p)-\underline{\alpha}(\omega)\right).\quad$ (Optimistic)
These formulations explicitly capture the robustness and flexibility required when handling stochastic thresholds. For solving each problem, we have the following assumptions:
Assumption 2 (Slater condition under pessimistic thresholds).
There exists $q^{\pi_{1}}\in\mathcal{F}_{\textup{pes}}$ and $\xi\in\mathbb{R}_{>0}^{m}$ such that $G^{T}q^{\pi_{1}}\geq\overline{\alpha}_{i}+\xi_{i}$ for all constraint $i\in[m]$. Set the Slater gap $\rho:=(r^{T}q^{\pi^{\star}}(p)-r^{T}q^{\pi_{1}})/\min_{i\in[m]}\xi_{i}$.
Assumption 3 (Slater condition under optimistic thresholds).
There exists $q^{\pi_{2}}\in\mathcal{F}_{\textup{opt}}$ and $\xi^{\prime}\in\mathbb{R}_{>0}^{m}$ such that $G^{T}q^{\pi_{2}}\geq\underline{\alpha}_{i}+\xi^{\prime}_{i}$ for all constraint $i\in[m]$. Set the Slater gap $\rho^{\prime}:=(r^{T}q^{\pi^{\star}}(p)-r^{T}q^{\pi_{2}})/\min_{i\in[m]}\xi_{i}^{\prime}$.
@@
Conditioned on Assumption 2, we have
$\min_{\lambda\in\mathbb{R}_{\geq 0}^{m}}\max_{q\in\mathcal{F}_{\textup{pes}}}\mathcal{L}(q,\lambda,\omega)=\max_{q\in\mathcal{F}_{\textup{pes}}}\min_{\lambda\in\mathbb{R}_{\geq 0}^{m}}\mathcal{L}(q,\lambda,\omega).$
@@
Under Assumption 2, we can regard Equation (1) as the following convex optimisation problem:
$q^{\pi^{\star}}\in\arg\max_{q^{\pi}\in\mathcal{F}_{\text{pes}}}r^{T}q^{\pi}\quad\text{s.t.}\quad Gq^{\pi}\geq\alpha,$ (6)
@@
    The objective function $f(\cdot)=-r^{T}(\cdot)$ is affine and therefore convex.
@@
    Each constraint $y_{i}(\cdot):=\alpha_{i}-g_{i}^{T}(\cdot)$ is affine and therefore convex.
@@
    By Assumption 2, Equation (6) admits at least one feasible solution, and since $\mathcal{F}_{\text{pes}}$ is compact while $f(\cdot)$ is continuous, an optimal solution exists.
@@
    A Slater point exists by Assumption 2, namely $q^{\pi_{1}}$.
@@
    All dual problems admit an optimal solution because the domain is compact and $f(\cdot)+\lambda^{T}y(\cdot)$ is continuous.
@@
Conditioned on Assumption 3, we have
$\min_{\lambda\in\mathbb{R}_{\geq 0}^{m}}\max_{q\in\mathcal{F}_{\textup{opt}}}\mathcal{L}(q,\lambda,\omega)=\max_{q\in\mathcal{F}_{\textup{opt}}}\min_{\lambda\in\mathbb{R}_{\geq 0}^{m}}\mathcal{L}(q,\lambda,\omega).$

### v3

@@ 4.1 Reward Regret and Constraint Violation of a Pessimistic Policy
Combining Step 2 and Step 3, and after rearrangement, we have
$\mathcal{L}(\lambda)\lesssim$ $\left(\rho+\frac{\|\lambda\|_{2}^{2}}{\rho}\right)\sqrt{mH^{2}T}+\left(\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}\right)$
$+\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA,$
Please note that this is an upper bound for the combination of reward and constraint violation. We can obtain the regret bound for reward or constraint violation through setting $\lambda=0$ or $\lambda_{i}=\rho e_{i}\mathbf{1}_{([\sum_{k=1}^{T^{\prime}}g_{i,k}]_{+}\neq 0)}$.
4.2 Reward Regret and Constraint Violations of an Optimistic Policy
When a certain degree of risk is acceptable in exchange for faster learning and higher reward potential (e.g., exploration‑heavy tasks or loosely constrained resource allocation), an optimistic threshold could be adopted in Algorithm 2, by setting $\alpha_{\diamond}^{t}=\underline{\alpha}^{t}$ at every episode $t$. Similarly, we prove reward regret and constraint violations of an optimistic policy as below.
@@ 4.2 Reward Regret and Constraint Violations of an Optimistic Policy
Under the same setting of Theorem 2 and Assumption 1, with probability $1-\delta$, Algorithm 2 under optimistic thresholds achieves the following regret bounds:
@@
$+\tilde{\mathcal{O}}\left(\sqrt{mSAH^{2}T/\gamma}+\sqrt{m}SAH\right).$
This theorem shows that, even under this aggressive stance, both reward regret and constraint violation remain sublinear over episodes $T$.
@@
The proof idea is similar with the pessimistic counterpart, so we discuss the most significant differences below due to space limitation. A detailed proof is given in Appendix H.
@@
We can derive a decomposition similar to Equation (4); however, unlike the pessimistic policy case, an additional term $\sqrt{\sum_{i=1}^{m}\left(\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right)^{2}}\|\lambda\|_{2}$ appears, which measures the total estimation error for all constraints. Applying Lemma 12, we have that with probability $1-\delta$,
$\sqrt{\sum_{i=1}^{m}\left(\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right)^{2}}\leq\tilde{\mathcal{O}}\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H+1)\sqrt{m}H^{2}SA+\sqrt{mSAH^{2}T/\gamma}\right).$
@@
Optimistic policies also exhibit difference, as below.
Lemma 5 (Policy optimality of optimistic policy).
Conditioning on the success event, for any episode $t\in[T^{\prime}]$, if we adopt the optimistic thresholds, then
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\leq\tilde{\mathcal{O}}\left((1+m\rho^{\prime})H^{2}\sqrt{T}\right).$
@@ 4.3 Safety-Performance Trade-Off
Our theoretical analysis presents a safety-performance trade-off: policies aiming for strict adherence to constraints can incur larger reward regret, while policies that aggressively maximise rewards are more prone to constraint violations. Concretely, Theorems 2 and 3 show that, compared to fixed-threshold baselines, pessimistic (strict) thresholds yield lower violations but higher reward regret, whereas optimistic thresholds exhibit the opposite behaviour.
@@
One may wonder whether it is possible to tune a hyperparameter to leverage this trade-off without resorting to two extreme cases – pessimistic or optimistic. To this end, we may introduce a hyperparameter $\xi\in[0,1]$ to interpolate between these extremes. In particular, instead of setting $\alpha_{\diamond}^{t}=\overline{\alpha}^{t}$ or $\underline{\alpha}^{t}$, we define a blended threshold
$\alpha_{\xi}^{t}=\xi\underline{\alpha}^{t}+(1-\xi)\overline{\alpha}^{t}$
at episode $t$. Varying $\xi$ adjusts how ‘strict’ or ‘relaxed’ the thresholds are, thereby controlling the balance between safety and performance. While our theorems currently focus on the two boundary cases $\xi=0$ (pessimistic) and $\xi=1$ (optimistic), it is straightforward to extend to an analogous analysis for intermediate $\xi$ values.
@@ 5 Conclusion
In this paper, we propose a novel primal-dual scheme for CMDPs with stochastic thresholds, which utilises samples from environmental interactions to estimate these thresholds. Building on this framework, we introduce SPOT, a primal-dual algorithm for episodic CMDPs with stochastic thresholds that is designed to balance regret minimisation with safety guarantees in an uncertain environment. Furthermore, we leverage either pessimistic or optimistic thresholds in SPOT, respectively. We theoretically prove that SPOT, under both threshold settings, achieves a regret bound of $\tilde{\mathcal{O}}(\sqrt{T})$ and a constraint violation of $\tilde{\mathcal{O}}(\sqrt{T})$ over $T$ episodes, which are in the same order as that of an approach with certain thresholds. Our findings reveal a subtle dance: pessimistic thresholds enforce stricter safety constraints that effectively reduce violations, albeit at the cost of increased reward regret; in the optimistic setting, a reverse situation is observed. This creates a ‘margin’ between safety and performance without relying on prior threshold information. More fundamentally, our treatment of thresholds as stochastic variables opens up a new design space for developing more varied and sophisticated safety mechanisms.
@@ Appendix A Summary of Notation
Action space $\mathcal{A}$, with cardinality AA
@@
Slater gap of $\pi^{0}$ $\rho\;=\;(r^{\top}q^{\pi^{\star}}(p)-r^{\top}q^{\pi^{0}})/\min_{i\in[m]}\xi_{i}$
@@
Threshold $\tilde{\alpha}_{i,h}(s,a)$ with $\mathbb{E}[\tilde{\alpha}_{i,h}(s,a)]=\alpha_{i,h}$
Constraint functions $y_{t}(s,a)=\alpha^{t}-Gq^{\pi_{t}}(p)$
@@
Occupancy measure (matrix form) $q=[q_{h}^{\pi}(s,a;p)]_{H\times S\times A}$
Lagrangian under stochastic thresholds $\mathcal{L}(q,\lambda):=r^{\top}q+\lambda^{\top}(G^{\top}q-\alpha)$
Lagrangian of Algorithm 2 $\mathcal{L^{\prime}}(q,\lambda):=\overline{r}^{\top}q+\lambda^{\top}(\overline{G}^{\top}q-\alpha_{\diamond})$
Optimal policy $q^{\pi^{\star}}\in\arg\max_{q\in\mathcal{F}_{q}}\min_{\lambda\in\mathbb{R}_{\geq 0}^{m}}\mathcal{L}(q,\lambda)$
@@
Step size $\eta_{\lambda}>0$ (hyperparameter)
@@
Visitation counter $N_{t,h}(s,a)=\sum_{l=1}^{t}\mathbf{1}\{s_{h}^{l}=s,a_{h}^{l}=a\}$
@@
Optimistic estimates $\overline{r}_{t},\;\overline{g}_{t},\;\overline{\alpha}_{t}$
@@ B.1 Lagrangian Formulation of CMDPs without Stochastic Thresholds
Under the bandit feedback paradigm, traditional CMDPs usually assume that rewards and constraints are stochastic or adversarial. Here, we take the stochastic case as an example, that is, rewards and constraints are randomly drawn from some probability distributions $\mathcal{R}$ and $\mathcal{G}_{i}$ for all $i\in[m]$. The immediate reward after taking an action $a$ at state $s$ is a random variable $\tilde{r}_{h}(s,a)\in[0,1]$ with expectation $\mathbb{E}_{\mathcal{R}}[\tilde{r}_{h}(s,a)]=r_{h}(s,a)$. Similarly, for every constraint $i\in[m]$, the immediate constraint $\tilde{g}_{i,h}(s,a)\in[0,1]$ for every state-action pair $(s,a)$ is a random variable with expectation $\mathbb{E}_{\mathcal{G}_{i}}[\tilde{g}_{i,h}(s,a)]=g_{i,h}(s,a)$.
@@
$\max_{q\in\Delta(\mathcal{M})}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}(q,\lambda):=r^{\top}q+\lambda^{\top}(G^{\top}q-\alpha),$
@@
In order to further discuss its strong duality, we make the following assumption, which is commonly used in the study of CMDPs [altman1999constrained, efroni2020exploration, li2024faster, ying2022dual, ding2022convergence, ding2023provably, muller2024truly, paternain2022safe].
Assumption 1 (Slater condition).
There exists $\pi^{0}\in\Pi$ and $\xi\in\mathbb{R}_{>0}^{m}$ such that $V_{g_{i}}^{\pi^{0}}\geq\alpha_{i}+\xi_{i}$ for all $i\in[m]$. Set the Slater gap $\rho=(r^{\top}q^{\pi^{\star}}(p)-r^{\top}q^{\pi^{0}})/\min_{i\in[m]}\xi_{i}$.
@@
Lemma 6 (Strong duality under fixed thresholds [paternain2019constrained]).
@@ B.2 Lagrangian Formulation of CMDPs with Stochastic Thresholds
From these definitions, it follows that the pessimistic and optimistic scenarios define the boundaries for all feasible policies, yielding: $\mathcal{F}_{\textup{pes}}\subseteq\mathcal{F}_{q}\subseteq\mathcal{F}_{\textup{opt}}$ with probability $1-\delta$ due to Theorem 1.
Our theoretical analysis follows Assumption 1. Building on this standard assumption, we further require that the pessimistic tightening of the thresholds does not eliminate the existence of such a strictly feasible point. Similar “conservative tightening” assumptions have appeared in recent studies [bura2022dope].
Assumption 2 (Slater condition).
For $i\in[m]$, we have $\overline{\alpha}_{i}-\alpha_{i}<\xi_{i}$.
@@
Under Assumptions 1 and 2, we have $\min_{\lambda}\max_{q\in\mathcal{F}_{\textup{pes}}}\mathcal{L}(q,\lambda)=\max_{q\in\mathcal{F}_{\textup{pes}}}\min_{\lambda}\mathcal{L}(q,\lambda).$
@@
Under Assumption Assumptions 1 and 2, we solve the following convex optimisation problem:
$q^{\pi^{\star}}\in\arg\max_{q\in\mathcal{F}_{\text{pes}}}r^{\top}q\quad\text{s.t.}\quad G^{\top}q\geq\overline{\alpha},$ (5)
@@
    The objective function $f(\cdot)=-r^{\top}(\cdot)$ is affine and therefore convex.
@@
    Each constraint $y_{i}(\cdot):=\overline{\alpha}_{i}-g_{i}^{\top}(\cdot)$ is affine and therefore convex.
@@
    Equation (5) admits at least one feasible solution, and since $\mathcal{F}_{\text{pes}}$ is compact while $f(\cdot)$ is continuous, an optimal solution exists.
@@
    A Slater point exists.
@@
    All dual problems admit an optimal solution because the domain is compact and $f(\cdot)+\lambda^{\top}y(\cdot)$ is continuous.
@@
Under Assumption 1, we have $\min_{\lambda}\max_{q\in\mathcal{F}_{\textup{pes}}}\mathcal{L}(q,\lambda)=\max_{q\in\mathcal{F}_{\textup{pes}}}\min_{\lambda}\mathcal{L}(q,\lambda).$